A slow month as eclipse prep and travel dominated most people’s time.
A few trips at the end of the month to check and characterize telescope operation. Dec oscillations on the 24” are plotted at left. Bob Harwell subsequently installed stops on the worm to restrict its lateral motion. Future tests will determine its effectiveness. (The plotted RA drift is less than an arcsecond/minute.)
An issue was found with SiTech pointing when absolute encoder values deviated too far from the brushless servo motor encoder values. The deviation was built up by manual operation of the mount, e.g. when balancing. A process is in place to ensure these values are sync’d on startup.
The filter wheel for use on the flipped optical path with the QHY290 camera is still on back order. Just before Labor Day, the PL16803 was sent FLI to have the liquid cooling system replaced with an air cooled system. Liquid cooling was too difficult to maintain. So neither optical path is available, at least till the end of September.